OpenStudy (anonymous):

Divide. (8m^7 – 10m^5) ÷ 2m^3

OpenStudy (callisto):

(8m^7 – 10m^5) ÷ 2m^3 = 8m^7 (÷ 2m^3) – 10m^5 (÷ 2m^3) = (8/2) m^(7-3) - (10/2) m^(5-3) = 4m^4 - 5 m^2